OpenStudy (anonymous):

a silver ball weighing 50g is removed from a furnace and dropped into a 350g of water at 0 degree celcius.if the equilibrium temperature is 10 degree celcius and specific heat capacity of platinum is 230 Jkg-1 oC-1, what is the temperature of the furnace?neglect the effect of the mass of the calorimeter.

OpenStudy (anonymous):

Silver ball or platinum? I am assuming that whatever it is, the ball is made of the material whose heat capacity is 230 JKg-1 oC-1

OpenStudy (anonymous):

Energy exchanged = m c (T2 - T1) Ball: 50/1000 X 230 X (x - 10) Water: 350/1000 X 4186 X (0 - 10) Total energy exchanged = 0 This means \[ \frac{50}{1000}230\left(x - 10\right) + \frac{350}{1000}4186\left(-10\right) = 0 \] Solve for x